OpenStudy (anonymous):

Find the equation of each line in slope intercept form: THrough the point (6,1) with slope = 1/3

OpenStudy (anonymous):

I got y=1/9x+1/3

OpenStudy (anonymous):

thats not quite right. you want it to be in the form y=mx+b where m is the slope and b is the y intercept. So if the slope is 1/3 you then have y=1/3x+b . You know that there is a point at (6,1) [remember that points are in the form (x,y)] we can know that there is a y intercept at 1, so your final equation is y=1/3x+1 hope that helps

OpenStudy (anonymous):

will it be y=1/3x-1 instead

OpenStudy (anonymous):

it would only be y=1/3x-1 if the point was (6,-1) not (6,1) which is what you put